# Who Said What? An Automated Approach to Analyzing Speech in Preschool Classrooms

### 05 Mar 2024 v2

@@ IV-C7 Lexical Overlap
Finally, we calculated the proportion of responses that contained none of the words in the preceding utterance (0 alignment). A high proportion of both teacher (Whisper=.65.65.65.65, expert=.74.74.74.74) and child responses (Whisper=.79.79.79.79, expert=.85.85.85.85) had no lexical overlap with the preceding utterance. Proportions of responses without lexical overlap appeared larger for the expert. This may be because Whisper relies on linguistic context in its transcriptions, and may be more likely to transcribe an utterance as containing words that it also detected in preceding and following utterances than is the expert.
IV-C8 Intraclass Correlations
@@
We also conducted intraclass correlations using audio files as the unit of analysis (see Table. III). Each audio file reflects the language environment of a given teacher or child. In reliability theory parlance, audio files are the objects of measurement and variance between audio files is true variance. Intraclass correlations ranged between .75.75.75.75 and .97.97.97.97. These intraclass correlations reflect audio file variance as a proportion of total variance. Total variance contains two sources of error–variance due to differences between Whisper and expert, and error variance. The intraclass correlations suggest that Whisper’s reliability in measuring child speech features was high (>.$>.83$) for all features with the exception of the rate of child questions and the proportion of teacher questions to which children responded (.75.75.75.75 and .74.74.74.74, respectively). Whisper’s reliability in measuring teacher speech features was high (>.$>.90$) for all features with the exception of the proportion of teacher responses to non-questions (.76.76.76.76).
@@ V Discussion
Preschool is a significant developmental context for the majority of 3- to 5-year olds in the United States. Within preschools, there is pervasive evidence that teacher speech features including question use and mean length of utterance are associated with children’s language outcomes [33, 34]. Likewise, children’s own speech production–including the phonemic complexity of their utterances– may be the bridge between teacher speech complexity and child language outcomes [11]. The language interactions that occur in preschool settings may have an outsize impact on children whose families live in poverty as well as children with developmental disabilities.
@@
Manual speaker classification and transcription is the limiting factor in understanding classroom speech. For example, expert transcription of the 110 minutes of audio data reported here took approximately 55 hours (5 hours of expert transcription per 10 minutes of audio). Researchers have begun to tackle this problem d through automated quantification of selected speech features from classroom audio [21, 22, 20]. We substantially add to this literature with an automated framework for the large-scale analysis of speaker classification and transcription (who said what).
@@
Speaker classification results indicated moderate levels of overall reliability for both teacher and child utterances. Word error rates–a rigorous metric–indicated relatively high levels of transcription accuracy for both teacher and child recordings. Comparison of expert and automated processing was not limited to standard reliability metrics but was extended to the speech features that are the likely substantive areas of analysis for classroom interaction research. These analyses used all available audio from both teacher and child recorders. The results suggest promising levels of correspondence on teacher and child MLU, rate of speech, use of questions, and responses to questions. However, each of these features must be examined individually. For example, our data suggest that Whisper over-estimated lexical alignment between utterances and their responses, likely based on the tendency of this large language model to “hear” words in responses that had been identified in adjacent utterances.
@@
Although descriptions of speech features are preliminary, both the automated framework and expert coding suggest that teachers have somewhat higher MLU than children and produce many more utterances than children. Teachers asked more questions than children but both teachers and children responded to about one-third of their interlocutor’s questions. The proportion of responses that contained words that were also contained in the preceding utterance was low. It should be noted, however, that interchanges such as those in Figure 4 (teacher: “What do you think is going to happen?” child: “it’s going to explode”) contain no lexical overlap. More sophisticated analyses of alignment–including semantic alignment based on meaning–are required [31].
@@
The current research has multiple limitations. Speaker classification accuracy requires improvement. Machine learning of Whisper transcriptions to bolster classification is one promising direction of research. Transcription results were relatively accurate but were limited to the teacher or child wearing the recorder, suggesting that the distance between mouth and recorder microphone might be a crucial variable. However, speech feature results showed impressive correspondence between automated and expert measurement both at an utterance level and an audio recording level.
@@ VI Conclusion
The current results suggest a framework for a breakthrough in the automated analysis of classroom speech. High-quality recorders worn by teachers and children yielded reliable automated transcription of what individuals said. Moreover, there were high levels of agreement in automated versus expert measurements of key features of their speech including MLU and question-asking. We have applied the automated pipeline to a dataset of classroom recordings from 13 children and 4 teachers observed on 17 occasions over one year. This dataset contains 765 hours of recordings and required 200 hours of processing time with the current automated pipeline. The automated pipeline yielded over half a million transcribed utterances, which we are currently analyzing. Thus, automated methods show great potential for producing datasets with which to understand interaction and development in naturalistic contexts.
@@ Acknowledgement
This work was supported by the National Institute on Deafness and Communication Disorders (R01DC018542) and the National Science Foundation (2150830).

### 10 Apr 2024 v3

@@ IV-C7 Lexical Alignment
Finally, we calculated the proportion of responses that contained none of the words in the preceding utterance (00 alignment). A high proportion of both teacher (Whisper=.78.78.78.78, expert=.88.88.88.88) and child responses (Whisper=.67.67.67.67, expert=.70.70.70.70) had no lexical overlap with the preceding utterance. All words in a pair of utterances are used to calculate lexical overlap. Proportions of responses without lexical overlap for child responses appeared larger for the expert than for Whisper. This may be because Whisper relies on linguistic context in its transcriptions, and may be more likely to transcribe a child utterance as containing words that it also detected in the preceding adult utterance than is the expert (increasing lexical overlap and decreasing the proportion of child responses with no lexical overlap). By contrast, there was a close match between expert and Whisper on the absence of lexical alignment in teacher responses to children.
IV-C8 Intraclass Correlation Approach
@@
For more comprehensive reliability evaluations, we conducted absolute level intraclass correlations using audio files as the unit of analysis (see Table III). Each audio file reflects the language environment of a given teacher or child. In reliability theory parlance, audio files are the objects of measurement and variance between audio files is true variance. These intraclass correlations reflect audio file variance as a proportion of total variance. Total variance contains two sources of error–variance due to differences in measurement (Whisper versus expert), and error variance.
IV-C9 Intraclass Correlation Results
Overall, intraclass correlations ranged between .74.74.74.74 and .97.97.97.97. Intraclass correlations for the rate of child question asking, the proportion of child questions that were responded to, and the proportion of teacher non-questions that were responded to by the child ranged from .$74-.76$, indicating only moderate reliability on these features. All other intraclass correlations–including child and teacher MLU and rate of question asking–were high (.$84-.97$). The results indicate that a substantial quantity of the variance in measured speech features was associated with differences in those features between audio clips.
@@ V Discussion
Preschool is an important developmental context for the majority of 3- to 5-year olds in the United States. Within preschools, there is pervasive evidence that teacher speech features including question use and MLU are associated with children’s language outcomes [37, 38]. Likewise, children’s own speech production may be a bridge between teacher speech complexity and child language outcomes [11]. Finally, language interactions that occur in preschool settings may have an outsize impact on children whose families live in poverty as well as children with developmental disabilities.
@@
Manual speaker classification and transcription is often a limiting factor in understanding classroom speech. For example, expert transcription of the $110$ minutes of audio data reported here took approximately $55$ hours ($5$ hours of expert transcription per 10 minutes of audio). Researchers have begun to tackle this problem through automated quantification of selected speech features from classroom audio [21, 22, 20]. We add to this literature with an automated framework for the large-scale analysis of speaker classification and transcription (who said what).
@@
Speaker classification results indicated moderate levels of overall reliability for both teacher and child utterances. Word error rates–a rigorous metric–indicated relatively high levels of transcription accuracy for both teacher and child recordings. Comparison of expert and automated processing was not limited to standard reliability metrics such as word error rate, but was extended to the speech features that are the likely substantive areas of analysis for classroom interaction research. These analyses used all available audio from both teacher and child recorders. The results suggest promising levels of correspondence on teacher and child MLU, rate of speech, use of questions, and responses to questions. However, each of these features must be examined individually. For example, our data suggest that Whisper over-estimated lexical alignment between teacher utterances followed by child utterances. This may be due to the tendency of this large language model to “hear” words in child utterances that had been identified in the previous teacher utterances.
@@
Although descriptions of speech features are preliminary, both the automated framework and expert coding suggest that teachers have somewhat higher MLU than children and produce many more utterances than children. Teachers asked more questions than children but both teachers and children responded to about one-third of their interlocutor’s questions. The proportion of responses that contained words that were also contained in the preceding utterance was low. It should be noted, however, that interchanges such as those in Fig. 4 (teacher: “What do you think is going to happen?” child: “it’s going to explode”) contain no lexical overlap. More sophisticated analyses of alignment–including semantic alignment based on meaning–are required [33].
@@
The current research has multiple limitations. Speaker classification accuracy requires improvement and current results were based on expert speaker classification. Machine learning of Whisper transcriptions to bolster speaker classification is one promising direction of research. Transcription results were relatively accurate but were limited to the teacher or child wearing the recorder, suggesting that the distance between mouth and recorder microphone might be a crucial variable. However, speech feature results showed impressive correspondence between automated and expert measurement both at an utterance level and an audio recording level.
@@ VI Conclusion
The current results suggest a framework for advancing the automated analysis of classroom speech. High-quality recorders worn by teachers and children yielded reliable automated transcription of what individuals said. Moreover, there were high levels of agreement in automated versus expert measurements of key features of their speech including MLU and question-asking. We have applied the automated pipeline to a dataset of classroom recordings from $13$ children and $3$ teachers observed on $17$ occasions over one year. This dataset contains $765$ hours of recordings and required $200$ hours of processing time with the current automated pipeline. The automated pipeline yielded over half a million transcribed utterances, which we are currently analyzing. Thus, automated methods show great potential for producing datasets with which to understand interaction and development in naturalistic contexts.
@@ Acknowledgement
This work was supported by the National Institute on Deafness and Communication Disorders (R01DC018542), the National Science Foundation (2150830), and the Simons Foundation Autism Research Initiative (SFI-AR-HUMAN-00004115-01).